Q- A particle moves in a region having a uniform magnetic field and parallel                   electric field.At some instant the velocity of a particle is perpendicular to the field     direction.The path of a particle will be?


    a) straight line


    b)circle


    c) helix with uniform pitch


    d) helix with non uniform pitch   EXPLAI?N ALS?O??

Grade:Upto college level

1 Answers

Chetan Mandayam Nayakar
13 years ago

At the instant of starting, the path is a circle.The electric field causes the particle to accelerate perpendicular to the plane of the circle caused by the magnetic field. Therefore the path is a helix

mv2/r = evB, r= mv/eB = Et/eB

In addition,because v increases with time, the helix has the shape of a stretched out, spring whose radius increases continuously from turn to turn. Therefore, the correct answer is option 'd'

'Win exciting gifts by answering the questions on Discussion Forum.'